# On the Variance of Neural Network Training with respect to Test Sets and Distributions

### v3

@@ Appendix B Proofs
$\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}(h))$ $=\mathop{\mathbb{E}}_{h\sim\mathcal{H}_{\mathcal{A}}}\big{[}\left(\mathrm{err}(h)-\mathbb{E}_{h}[\mathrm{err}(h)]\right)^{2}\big{]}$
$=\mathop{\mathbb{E}}_{h}\left[\left(\mathop{\mathbb{E}}_{x,y}[\mathrm{err}_{x,y}(h)]-\mathop{\mathbb{E}}_{h}[\mathop{\mathbb{E}}_{x,y}[\mathrm{err}_{x,y}(h)]]\right)^{2}\right]$
$=\mathop{\mathbb{E}}_{h}\left[\left(\mathop{\mathbb{E}}_{x,y}[\mathrm{err}_{x,y}(h)]-\mathop{\mathbb{E}}_{x,y}[\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x,y}(h)]]\right)^{2}\right]$
$=\mathop{\mathbb{E}}_{h}\left[\left(\mathop{\mathbb{E}}_{x,y}[\mathrm{err}_{x,y}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x,y}(h)]]\right)^{2}\right]$
$=\mathop{\mathbb{E}}_{h}\left[\mathop{\mathbb{E}}_{x_{1},y_{1}}[\mathrm{err}_{x_{1},y_{1}}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x_{1},y_{1}}(h)]]\cdot\mathop{\mathbb{E}}_{x_{2},y_{2}}[\mathrm{err}_{x_{2},y_{2}}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x_{2},y_{2}}(h)]]\right]$
$=\mathop{\mathbb{E}}_{h}\mathop{\mathbb{E}}_{x_{1},y_{1}}\mathop{\mathbb{E}}_{x_{2},y_{2}}\left[(\mathrm{err}_{x,y}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x,y}(h)])(\mathrm{err}_{x_{2},y_{2}}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x_{2},y_{2}}(h)])\right]$
$=\mathop{\mathbb{E}}_{x_{1},y_{1}}\mathop{\mathbb{E}}_{x_{2},y_{2}}\mathop{\mathbb{E}}_{h}\left[(\mathrm{err}_{x_{1},y_{1}}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x_{1},y_{1}}(h)])(\mathrm{err}_{x_{2},y_{2}}(h)-\mathop{\mathbb{E}}_{h}[\mathrm{err}_{x_{2},y_{2}}(h)])\right]$
@@
$=\,\,(1/n)\cdot\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}}\left[\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}_{x,y}(h))\right]\,+\,(1-1/n)\cdot\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}(h))$
@@ B.2 Theorem 2
The quantity $\hat{\sigma}_{S}^{2}$ is also equal to $\binom{n}{2}^{-1}\sum_{i=1}^{n}\sum_{j\neq i}\mathop{\mathrm{Cov}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}_{x_{i},y_{i}}(h),\mathrm{err}_{x_{j},y_{j}}(h))$. Comparing this formula to Lemma 1 may help provide intuition for why it is an estimator for the distribution-wise variance. The formulation given in Theorem 2 looks less intuitive, but the benefit is that we only have to calculate $n$ separate variances, rather than $\binom{n}{2}$ separate covariance values.
We note that the proofs of Theorem 1 and Theorem 2 do not assume anything about the error function $\mathrm{err}_{x,y}(h)$ [^2] That is, other than it being non-pathological enough to allow the interchanges of expectation via Fubini’s theorem, e.g., nonnegative suffices., so, e.g., they are also true for regression tasks.
@@ B.3 Theorem 3
Define the random variable $q(x)=\mathop{\mathbb{E}}_{h\sim\mathcal{H}_{\mathcal{A}}}[1\{h(x)=1\}]$ to be the proportion of training runs which classify $x$ as positive, with the randomness being over $x\sim\mathcal{D}$. We first compute a formula for $\mathrm{err}(\mathcal{A})$ in terms of $q$. By the usual laws of conditional expectation we have
@@
Using the assumption of class-wise calibration, this is equal to
$\mathop{\mathbb{E}}_{q}[q(1-q)+(1-q)q]=\mathop{\mathbb{E}}_{q}[2q(1-q)].$
Next we analyze the example-wise variance. We have
@@
The second equality uses the formula for variance of a Bernoulli variable. The third equality uses the fact that, regardless of whether $y=0$ or $y=1$, the product $\mathop{\mathbb{E}}_{h}[1\{h(x)\neq y\}](1-\mathop{\mathbb{E}}_{h}[1\{h(x)\neq y\}])$ is equal to $\mathop{\mathbb{E}}_{h}[1\{h(x)=1\}]\mathop{\mathbb{E}}_{h}[1\{h(x)=0\}]$. The fourth equality applies the assumption of class-wise calibration.
Finally by combining Lemma 2 with the previous two results we get
@@
$=(1-1/n)\cdot\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}(h))\,+\,(1/n)\cdot\mathop{\mathbb{E}}_{q}[q(1-q)]$

### v4

@@ Appendix B Proofs
$\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}(h))$ $=\mathop{\mathbb{E}}_{h\sim\mathcal{H}_{\mathcal{A}}}\big{[}\left(\mathrm{err}(h)-\mathbb{E}_{h^{\prime}}[\mathrm{err}(h^{\prime})]\right)^{2}\big{]}$
$=\mathop{\mathbb{E}}_{h}\left[\left(\mathop{\mathbb{E}}_{x,y}[\mathrm{err}_{x,y}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x,y}(h^{\prime})]]\right)^{2}\right]$
$=\mathop{\mathbb{E}}_{h}\left[\mathop{\mathbb{E}}_{x_{1},y_{1}}[\mathrm{err}_{x_{1},y_{1}}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x_{1},y_{1}}(h^{\prime})]]\cdot\mathop{\mathbb{E}}_{x_{2},y_{2}}[\mathrm{err}_{x_{2},y_{2}}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x_{2},y_{2}}(h^{\prime})]]\right]$
$=\mathop{\mathbb{E}}_{h}\mathop{\mathbb{E}}_{x_{1},y_{1}}\mathop{\mathbb{E}}_{x_{2},y_{2}}\left[(\mathrm{err}_{x_{1},y_{1}}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x_{1},y_{1}}(h^{\prime})])(\mathrm{err}_{x_{2},y_{2}}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x_{2},y_{2}}(h^{\prime})])\right]$
$=\mathop{\mathbb{E}}_{x_{1},y_{1}}\mathop{\mathbb{E}}_{x_{2},y_{2}}\mathop{\mathbb{E}}_{h}\left[(\mathrm{err}_{x_{1},y_{1}}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x_{1},y_{1}}(h^{\prime})])(\mathrm{err}_{x_{2},y_{2}}(h)-\mathop{\mathbb{E}}_{h^{\prime}}[\mathrm{err}_{x_{2},y_{2}}(h^{\prime})])\right]$
@@
$=(1/n)\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}}\left[\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}_{x,y}(h))\right]\,+\,(1-1/n)\mathop{\mathbb{E}}_{(x_{1},y_{1}),(x_{2},y_{2})}\left[\mathop{\mathrm{Cov}}_{h}(\mathrm{err}_{x_{1},y_{1}}(h),\mathrm{err}_{x_{2},y_{2}}(h))\right]$
$=\,\,(1/n)\cdot\mathop{\mathbb{E}}_{(x,y)\sim\mathcal{D}}\left[\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}_{x,y}(h))\right]\,+\,(1-1/n)\cdot\mathop{\mathrm{Var}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}(h)).$
@@ B.2 Theorem 2
The quantity $\hat{\sigma}_{S}^{2}$ is also equal to $\binom{n}{2}^{-1}\sum_{i=1}^{n}\sum_{j\neq i}\mathop{\mathrm{Cov}}_{h\sim\mathcal{H}_{\mathcal{A}}}(\mathrm{err}_{x_{i},y_{i}}(h),\mathrm{err}_{x_{j},y_{j}}(h))$. Comparing this formula to Lemma 1 may help provide intuition for why it is an estimator for the distribution-wise variance. The formulation given in Theorem 2 looks less intuitive, but the benefit is that we only have to calculate $n$ separate variances, rather than $\binom{n}{2}$ covariances.
We note that the proofs of Theorem 1 and Theorem 2 do not assume anything about the error function $\mathrm{err}_{x,y}(h)$ [^2] Other than it being non-pathological enough to allow the interchanges of expectation via Fubini’s theorem., so, e.g., they are also true for regression tasks.
@@ B.3 Theorem 3
Define the random variable $q(x)=\mathop{\mathbb{E}}_{h\sim\mathcal{H}_{\mathcal{A}}}[1\{h(x)=1\}]$ to be the proportion of training runs which classify $x$ as positive, with the randomness being over $x\sim\mathcal{D}$. We first obtain a formula for $\mathrm{err}(\mathcal{A})$ in terms of $q$. By the usual laws of conditional expectation we have:
@@
Using the assumption of class-wise calibration, this formula simplifies to $\mathop{\mathbb{E}}_{q}[2q(1-q)]$. Next we analyze the example-wise variance. We have:
@@
Where the second equality uses the formula for variance of a Bernoulli variable. The third equality uses the fact that, regardless of whether $y=0$ or $y=1$, the product $\mathop{\mathbb{E}}_{h}[1\{h(x)\neq y\}](1-\mathop{\mathbb{E}}_{h}[1\{h(x)\neq y\}])$ is equal to $\mathop{\mathbb{E}}_{h}[1\{h(x)=1\}]\mathop{\mathbb{E}}_{h}[1\{h(x)=0\}]$. The fourth equality applies the assumption of class-wise calibration.
Combining the above two results yields $\mathop{\mathbb{E}}_{x,y}[\mathop{\mathrm{Var}}_{h}(\mathrm{err}_{x,y}(h))]=\mathrm{err}(\mathcal{A})/2$. Therefore by Lemma 2 we have:
@@ Appendix C ImageNet and distribution shifts
The above result provides empirical confirmation of a theory advanced by Cohen-Wang et al. (2024). The authors observed that pretraining provides a greater performance benefit for models evaluated on out-of-support distribution shifts like ImageNet-Sketch than it does for in-support shifts like ImageNet-V2. Seeking to understand this phenomenon, they analyzed a simple logistic regression setting and found that without pretraining, out-of-support shifts can induce more variance in behavior between runs of training than in-support shifts, due to a greater dependence on the initialization. Our above result, namely that ImageNet-Sketch has a distribution-wise standard deviation of 0.257% across repeated ResNet-18 trainings whereas ImageNet-V2 has only 0.071% (13×13$\times$13 × less variance), provides evidence that the logistic regression-based theory of Cohen-Wang et al. (2024), beyond being a source of useful intuition, is also directly true for neural networks.